OpenStudy (anonymous):

family income 32,000. they spent 1//5 for housing. how much spent on housing

OpenStudy (anonymous):

\[\frac{ x }{ 32000 }=\frac{ 1 }{ 5}\] 5x=32000 x=32000/5 x=$6400

OpenStudy (anonymous):

you divide 1/5 and that comes to a decimal .2 so then you divide .2 into 32000 and come out with 160,000

OpenStudy (anonymous):

thank you. can you help me with 2/3 divided by 1/7

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } \div \frac{ 1 }{ 7 }\] Multiply the first by the reciprocal of the second \[\frac{ 2 }{ 3 }\times \frac{ 7 }{ 1 }\]=\[\frac{ 14 }{ 3 }\]

OpenStudy (doc.brown):

here's why:\[\dfrac{\dfrac{2}{3}}{\dfrac{1}{7}}\]\[\dfrac{\dfrac{2}{3}}{\dfrac{1}{7}}\times\dfrac{\dfrac{7}{1}}{\dfrac{7}{1}}\]